Systems Staff Meeting Notes 05/20/02

Systems Staff Meeting Notes
May 20, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Cynthia notes that Star Wars was excellent. It is her new favorite movie.
  2. Bill notes that it is a very long walk across campus to Traffic & Parking.
  3. Fire drill this morning – was not audible in the vault. It was also not detectable from the loading dock – Bill was coming in that way while the alarm was going off. The sheet from PAULA was not printed out. Now, as a rule, the first person listed “at desk” will be responsible for printing out the list.
Review Minutes of 5/13/02
  1. Ann #1 – Ying has returned to badminton!
  2. Equip RIR #2 – Trend anti-virus for the server – has become a hassle trying to include it under our current contract. Trend has lost any record of the previous transaction.
  3. Equip RIR #3 – HP 2200D printers – old disk found to install drivers. Disk has been duplicated, distributed and the printers are now functioning.
  4. Nets #3 – SAN fail-over has been fixed, no need to move profiles.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Dell – secondary backup server for MSL – has the order gone out? Unsure, need to double-check.
  2. Order for 30 Dells has gone out.
  3. Wiring for MSL is out – Tuesday, May 21 – wiring to take place.
  4. Four switches for the MSL project were received.
  5. 15 UPS received – swapped for ILLIAD and Squid.
  6. All of the optical mice have been distributed.
  7. Need to schedule down time for installation of UPS’s. Early in the AM in the upcoming weeks will be good.
  8. Problematic workstation: David Hickey’s – it was a W2K upgrade. It has bit-rot and is especially problematic because of all the foreign language software he has on it. There is also a liaison jurisdiction issue between MSL & Collection Management in this case. We need to make sure that he is completely backed up on the server then reinstall W2K from scratch, with a liaison involved.
Liaisons
  1. Classroom workstation – all have been physically installed. Eight machines were done as of Friday; all twelve should be set up by noon today. Everyone has signed off that the image is OK, even ‘perfect’ and ‘exactly what they wanted’. Classroom to be back in use by May 23 or 24.
  2. Front desk monitor configuration – Flat screens are not designed to swivel around to show the public what’s on the screen. We could add a separate flat screen that would be software-controlled to display or not display to the patron. There is a meeting scheduled today to discuss the issue and try to come to some resolution.
Networking
  1. The Kleez virus is more active than ever this week – 40 hit yesterday - all were caught by the scanner. There is an increase in the number of email impersonations – there is no known way to stop it. Kleez is reactivating old viruses like Majestra. All we can do is hope that the scanner continues to catch it.
  2. Backups showed that the Exchange server was stopped Friday at about 6PM.
  3. Disk quotas – user folders on NT2R set to 2 GIG. User will get a warning at 1.5 GIG. We need to rearrange the set-up on NT2R Department folders need to be at the top of SAN – Suzy will have some links to fix after the change. May have to redirect the virtual directory.
  4. No news on the MSL power.
  5. Security – stable, waiting for servers. According to the Dell site, they should ship out next Tuesday.
  6. Central defrag & software install – may not happen this year due to a budget hold.
Software Projects Status
  1. Current Net- finished, Bobbie is still evaluating.
  2. IRDB – small fix last week, it is on the list with the Trouble Tickets - to be made more “user-friendly”.
  3. Software/Hardware Requests online – SysRequest – is ongoing.
  4. Erich has a new list of changes for Bryant – it is now a full-fledged project.
  5. Sanborn – complicated endeavor – old fire insurance maps – Erich wants them all in the database. This will take awhile.
  6. ILL – delivery through ILLIAD – unknown – Winston made a request, got a message that it was submitted, but no further updates.
  7. Complications with LAD scanning on Ariel.
  8. Directory Services on campus – switching from SSN to UF ID. Scheduled to occur in December. ALEPH will not be ready. Patron record is currently SSN. Need to meet with FCLA & Warren Curry – this has potential to be a real challenge.
Web Projects
  1. Suzy hired Nithya to check links. She is currently working on the Personnel site, since they requested help with it.
  2. Problems with the Stat server- SPIT group was intrigued – they think the usability of the Web server is top priority. The top level seems to be choking on the reports – maybe too many archived files.
DP Coordinator
  1. Students in place and operational.


Back to the Systems Homepage

Last updated June 10, 2002
by Debra Harris
debharr@mail.uflib.ufl.edu